# AIMCoT: Active Information-driven Multimodal Chain-of-Thought for Vision-Language Reasoning

### v1

@@ 5.4 Ablation Study
The performance gains from AIMCoT are most pronounced on LLaVA-W, a challenging benchmark requiring open-ended generation. This large improvement starkly demonstrates our model’s advanced capability to comprehend intricate multimodal information and tackle demanding, unconstrained tasks.
@@
5.5 In-depth Analysis: the Interplay between CAG and AVP
Our proposed components collaborate organically to foster the construction of multimodal CoT: as a preceding module, CAG enriches the context to benefit the construction of the candidate set; subsequently, AVP triggered by DAT proactively selects the most salient regions from the candidate set.
In this section, we investigate the interaction between CAG and AVP via an ablation study. Starting from the baseline ICoT model (BM), i.e., AIMCoT stripped of all proposed modules, we sequentially add CAG and then AVP. Table 4 presents the results, from which we derive the following insights:
  • The compatibility between CAG and AVP: both CAG and AVP individually provide significant gains, but their combination synergistically improves the construction of the interleaved CoT.
  • The consistent superiority of AVP over Top-K selection: AVP consistently and significantly outperforms the standard Top-K selection method for choosing image regions, both with and without the presence of the CAG module.
  • The interplay between CAG and AVP: The average performance improvement of AVP over Top-K selection increases from 2.62% to 2.94% when CAG is introduced. This suggests CAG enhances the source attention map, providing a more reliable set of candidate regions and thereby unlocking AVP’s full potential to select the most salient visual evidence.
@@ 6 Conclusion
In this paper, we propose AIMCoT, a novel framework that reframes the construction of interleaved-modal CoT as an active, information-foraging process, addressing the limitations in existing methods, which often rely on passive, heuristic-driven mechanisms for selecting and inserting visual information at suboptimal moments. Our extensive experiments on three popular and challenging benchmarks demonstrate that AIMCoT significantly outperforms state-of-the-art methods in both 0- and 1-shot settings (up to 18%). By dynamically structuring its reasoning and actively seeking the most informative visual cues, AIMCoT achieves a more proactive, goal-oriented, and human-like approach to vision-language reasoning.
Despite the strong performance, AIMCoT presents avenues for future exploration. The AVP module, while highly effective and optimized, introduces a slight computational overhead compared to simpler attention-based selection. Future work could explore lightweight, learnable policies for region selection to further enhance its deployability. We also plan to extend our evaluation to a broader range of VLM architectures and more complex, long-form reasoning tasks to further probe the generalizability and limits of our active information-seeking paradigm.
7 Reproducibility Statement
To ensure the reproducibility of our work, we provide the complete source code for our AIMCoT framework and all experiments in the supplementary materials, accessible via an anonymized link: https://anonymous.4open.science/r/AIMCoT. The architectural details and theoretical underpinnings of our proposed components, including Context-enhanced Attention-map Generation (CAG), Active Visual Probing (AVP), and Dynamic Attention-shifting Trigger (DAT), are thoroughly described in Section 4 of the main paper. The specific greedy algorithm employed by AVP is detailed in Algorithm 1 (Appendix E). All datasets used in our evaluation are publicly available benchmarks, as detailed in Section 5.1 and Appendix C. We provide a comprehensive list of all hyper-parameter settings used to achieve the reported results for each benchmark in Appendix G.2 (Table 5). Furthermore, extensive ablation studies (Section 5.4), and in-depth analyses (Section 5.5, Appendices F,H, I, K, and L), are provided to allow for a complete replication of our findings.
Appendix A LLM Usage Statement
In the preparation of this manuscript, we utilized a large language model (LLM) as a general-purpose writing assistance tool. The primary uses of the LLM were twofold:
  • Grammatical Correction: The LLM was employed to proofread the manuscript for grammatical errors, spelling mistakes, and awkward phrasing. This helped improve the overall clarity and readability of our paper.
  • Text Condensation: To adhere to the page limits of the conference, we used the LLM to help condense and rephrase certain paragraphs and sentences.
All suggestions provided by the LLM were carefully reviewed, critically evaluated, and manually edited by the authors to ensure that the scientific integrity and intended meaning of our work were preserved. Notably, the LLM was not used for core research activities, including the ideation of the AIMCoT framework, the design of experiments, the analysis of results, or the generation of the primary scientific claims. The final intellectual content and all contributions presented in this paper are entirely our own.
Appendix B Explanation for Figure 1
Figure 1 shows the images of the 22nd question on LLaVA-W benchmark. The query for this question is: "What’s the name of the restaurant serving these dishes?" and the image is a close-up photo of a meal at ICHIRAN. The left figure is the original image, and the right figure visualizes the regions selected by the Top-K strategy. The num_selected_patchesnum\_selected\_patches regions with the highest, second-highest, and third-highest scores are designated as the first, second, and third sets, colored red, purple, and blue, respectively. The first set is utilized by the baseline model according to its default design (num_selected_patches=72num\_selected\_patches=72).
Appendix C Introduction to the Benchmarks
@@
M3CoT Chen et al. (2024) is a novel multimodal CoT benchmark, which introduces complex, multi-step problems across science, mathematics, and commonsense domains, comprising 11,459 samples in total. M3CoT is characterized by succinct textual queries (<15 tokens on average) paired with intricate problems. This inherent text-vision imbalance makes it an ideal platform to validate the efficacy of our proposed CAG in mitigating this issue and the superiority of AVP in proactively selecting the salient visual regions.
@@ Appendix D Introduction to the Baseline Models
Appendix E Greedy Algorithm within AVP Module
The complete process of the greedy algorithm within AVP is shown in Algorithm 1.
Input: total candidate set $C$, size of optimal selection $K$
Output: optimal selection $S$
1 $R^{*}\leftarrow\varnothing$
2 for $k\leftarrow 1,2,\cdots,K$ do
3 $U_{B}\leftarrow H(Y|I,x,y_{<t},R^{*})$
4 for $i\leftarrow 1,2,\cdots,N+M$ do
5    $U_{C,i}\leftarrow H(Y|I,x,y_{<t},R^{*}\cup\{R_{i}\})$
6    $IG(\{R_{i}\})\leftarrow U_{B}-U_{C,i}$
7 $R_{next}\leftarrow\text{argmax}_{R_{i}\in C\backslash R^{*}}\{IG(\{R_{i}\})\}$
8 $R^{*}\leftarrow R^{*}\cup\{R_{next}\}$
9
[…]
11 return $S$
Algorithm 1 Greedy Algorithm for Optimal Region Selection
Appendix F Detailed Analysis of Key Moments to Insert Visual Information
Formal Definition of Attention Shifts. To analyze attention shifts, we examine the averaged attention maps across all attention heads in the last three layers of the VLM during the prediction of each token $t$. The model’s total attention scores allocated to the visual and text components of the input are respectively measured as follows:
$A_{visual}(t)=\sum_{i\in\text{indices of $C_{visual}$}}\bar{a}_{t,i},\hskip 14.22636ptA_{text}(t)=\sum_{j\in\text{indices of $C_{text}$}}\bar{a}_{t,j},$ (12)
where $C_{visual},C_{text}$ are the visual and text information within the context, respectively. Then, the shift in attention from the textual to the visual modality while generating token $t$ is defined as follows:
$\delta_{t}=A_{visual}(t)-A_{visual}(t-1).$ (13)
Based on these concepts, we design a two-part experiment:
To group the generated predictions according to response quality, we establish high- and low-scoring groups. All predictions are ranked in descending order by their ROUGE-L scores. The top 30% form $G_{h}$, the high-scoring group (high-quality responses), and $G_{l}$, the bottom 30% form the low-scoring group (low-quality responses).
Appendix G Model Settings
G.1 Template of $\mathcal{P}_{CAG}$
Figure 4 provides an intuitive example showing the template of $\mathcal{P}_{CAG}$ and how it is used to prompt the VLM to carefully generate a guiding description for the input image. In particular, for multiple-choice questions, such as those in M3CoT and ScienceQA, we prepend the following brief explanation to $\mathcal{P}_{CAG}$ to aid the VLM in better understanding its designated task: "This is a multiple-choice question. The question is based on the image provided."
@@ G.1 Template of $\mathcal{P}_{CAG}$
G.2 Setting of Hyper-parameters
The hyper-parameter settings are presented in Table 5. For easier understanding, we provide further explanation of the hyper-parameters $s_{r}$ and $s_{g}$ as follows: in the AVP module, the input image is first divided into $s_{g}\times s_{g}$ grids according to the set "Grid size" $s_{g}$. Then, each region finally selected by AVP consists of $s_{r}\times s_{r}$ grids.

### v2

@@ 5.1 Evaluation Setup
5.2 Performance Comparison
Table 1 reports the experimental results. AIM-CoT consistently outperforms all baselines (both text-only and I-MCoT) across all benchmarks and backbones, under both 0- and 1-shot settings. This demonstrates the strong advantage of AIM-CoT for complex vision-language reasoning (besides, the analyses in Appendix M show that AIM-CoT maintains good deployability).
More specifically, both I-MCoT methods (AIM-CoT and ICoT) surpass the text-only baselines, confirming the efficacy of explicit incorporation of visual evidence. Moreover, although AIM-CoT and ICoT follow the same I-MCoT paradigm, AIM-CoT achieves stronger performance. This gain comes from three key improvements: (1) AIM-CoT uses CAG to directly refine the VLM’s cross-attention distribution, instead of grounding in the unreliable signal used by ICoT; (2) it replaces fragile attention-based heuristics with information-driven selection (AVP); and (3) it replaces a static trigger with a dynamic and more precise trigger (DAT).
Finally, the performance gains vary across backbones. In Appendix K, we analyze this effect in detail and attribute it to the interaction between model architecture and model scale.
5.3 Ablation Study
In this section, we conduct the ablation study to verify the efficacy of each component within AIM-CoT. The detailed settings are as follows:
  • w/o CAG: The VLM operates under the image $I$ and the raw query $x$, instead of the CAG-enhanced query $x^{\prime}$.
  • w/o AVP: The information-driven selection (AVP) is replaced by attention-based selection.
  • w/o DAT: DAT is substituted with a static trigger (i.e., visual evidence is inserted whenever the VLM outputs a newline character).
The results on Chameleon-7B are shown in Table 2 (more results and analyses on other backbones are deferred to Appendix J). First, refining the VLM’s attention distribution via CAG can strongly improve VLM’s performance on VQA tasks. This is because the attention is not only the basis of model generation but also one source of candidate evidence. Furthermore, (1) even without CAG, AIM-CoT still significantly outperforms ICoT (Table 1); and (2) removing AVP and DAT causes a larger performance drop than removing CAG. These results highlight the importance of tackling the what to see and when to see it questions. In response to these two issues, we propose AVP and DAT, respectively.
To provide a comprehensive understanding of AIM-CoT, we go beyond merely validating component efficacy. In Sections 5.55.6, and Appendix H, we present both quantitative and visualized qualitative analyses of AVP, elucidating why information-driven regions are more effective than attention-based regions.
5.4 Interplay between CAG and AVP
In this section, we investigate the interaction between CAG and AVP. This is achieved by sequentially adding them to a basic model (BM) (i.e., AIM-CoT stripped of all its components).
As shown in Table 3, while CAG and AVP yield individual gains, their combination exhibits a clear phenomenon of super-additivity, i.e., the joint improvement consistently exceeds the sum of their separate contributions. Specifically, on M3CoT, the joint gain ($1.0\%$) significantly surpasses the linear sum of individual gains ($0.3\%+0.4\%=0.7\%$). This trend holds across all benchmarks (e.g., $1.7\%>1.4\%$ on ScienceQA).
This non-linear boost validates the interplay between the modules: CAG refines the VLM’s attention distribution, which amplifies AVP’s effectiveness; the AVP, in turn, precisely mines the salient visual evidence from the attention-driven candidates ($C_{attn}$) enhanced by CAG.
@@ 5.4 Interplay between CAG and AVP
5.5 Quantitative Analysis: Semantic Relevance
We further examine whether the visual regions selected by AIM-CoT match human-centered perception and the needs of the task. Our premise is that informative evidence should correspond to distinct and complete semantic entities rather than repetitive background patterns. We therefore use the Segment Anything Model (SAM) Kirillov et al. (2023) as a proxy to assess whether the selected regions align with meaningful object concepts.
We conduct this evaluation on M3CoT and ScienceQA, randomly sampling 500 instances from each benchmark. Results are averaged over the four backbones: Chameleon-7B, Janus-Pro-7B, Qwen2-VL-7B, and Qwen2.5-VL-32B. To build a reference for visually relevant evidence, we use SAM to generate segmentation masks for the key entities mentioned in the question and ground-truth answer. We then define the Semantic Hit Rate (SHR) as the percentage of instances in which the selected regions substantially overlap with the corresponding SAM masks, measured by IoU greater than 0.5.
Table 4: Comparison of SHR on M3CoT and ScienceQA (subset of 500 samples each). The results are averaged across four backbones.
Method Selection Strategy SHR $\uparrow$
M3CoT ScienceQA
ICoT Gao et al. (2025) Attention-driven Top-K 18.7% 24.3%
AIM-CoT (Ours) Info. Gain-driven AVP 65.2% 71.8%
Table 4 shows that the attention-driven Top-K strategy used by ICoT struggles to recover semantically relevant evidence, with SHR dropping to 18.7% on M3CoT. This supports the observation that raw attention maps often drift toward high-contrast background patterns or generic salient objects, rather than the specific evidence required for reasoning. In contrast, AIM-CoT achieves substantially higher SHR scores, reaching 65.2% on M3CoT and 71.8% on ScienceQA. These results indicate that our active information-seeking paradigm filters out distractors and selects regions that are both structurally coherent and semantically aligned with the question, making them closer to human judgments of relevant evidence.
5.6 Quantitative Analysis: Alignment with Human Intuition
We next evaluate whether the visual regions selected by AIM-CoT better align with human intuition than those selected by the attention-driven baseline. Following prior work, we use GPT-4v Achiam et al. (2023) as an expert proxy judge, as strong multimodal judges have been shown to correlate well with human preferences Huang et al. (2025); Gera et al. (2025); D’Souza et al. (2025); Findeis et al. (2025); Zhen et al. (2025).
We evaluate 500 randomly sampled instances from the M3CoT and LLaVA-W benchmarks. For each instance, GPT-4v is given the textual query, the original image, and two anonymized sets of selected regions: one produced by ICoT (Top-K) and the other by AIM-CoT (AVP). We randomize the order of the two sets to eliminate position bias. GPT-4v then compares them along three criteria: semantic relevance, object completeness, and overall helpfulness. The evaluation prompt is shown in Table 5.
Table 5: The core prompt used for GPT-4v blind pairwise comparison.
System Prompt: You are an expert judge. Compare two sets of image crops (Set A and Set B) extracted from the Original Image. Decide which set is more helpful for a human to answer the Question. Criteria: 1. Relevance: Does the set contain the specific objects or details mentioned in the question? 2. Completeness: Are the objects complete, or are they meaningless background noise/fragments? 3. Helpfulness: Which set would better help a human answer the question without seeing the full image? Output: ’Set A is better’, ’Set B is better’, or ’Tie’.
Table 6: GPT-4v preference rates on 500 samples from M3CoT and LLaVA-W. “Win” denotes that AIM-CoT provides better visual evidence than ICoT. “Tie” indicates equal quality.
Benchmark AIM-CoT Win Tie ICoT Win
M3CoT (Reasoning) 76.4% 16.2% 7.4%
LLaVA-W (In-the-Wild) 81.2% 13.6% 5.2%
As shown in Table 6, AIM-CoT is strongly preferred across both benchmarks. It wins on 76.4% of M3CoT samples and 81.2% of LLaVA-W samples, while only a small fraction favor ICoT. These results suggest that AIM-CoT more consistently identifies complete and semantically informative evidence that matches human intuition. GPT-4v’s qualitative judgments further indicate that ICoT often focuses on noisy regions, whereas AIM-CoT more reliably localizes the full objects needed for reasoning.
@@ 6 Conclusion
In this paper, we propose AIM-CoT, a novel I-MCoT framework that aims to frame the construction of interleaved-modal CoT as an active information-foraging process. Existing methods’ static triggers fail to capture the dynamic needs of the VLM for visual information in complex reasoning, and their attention-based selectors depend heavily on the VLM’s attention, which can be unreliable (for selection). In response to these challenges, AIM-CoT dynamically monitors the VLM’s cognitive need for fine-grained visual evidence for subsequent generation, and accordingly, selects salient visual evidence in an information-driven manner. Extensive experiments demonstrate that AIM-CoT outperforms the leading methods across three benchmarks and four VLM backbones.
7 Acknowledgements
This work was partially supported by the Guangdong Basic and Applied Basic Research Foundation under Grant No. 2024A1515011949 and 2026A1515011672, the Shenzhen Science and Technology Program under Grant No. GXWD20231130110308001 and JCYJ20250604145617023.
Appendix A Experimental Setup and Reproducibility
A.1 Benchmarks
@@
M3CoT Chen et al. (2024a) is a novel multimodal CoT benchmark, which introduces complex, multi-step problems across science, mathematics, and commonsense domains, comprising 11,459 samples in total. M3CoT is characterized by succinct textual queries (<15 tokens on average) paired with intricate problems. This inherent text-vision imbalance makes it an ideal platform to validate the efficacy of our proposed CAG in mitigating this issue and the superiority of AVP in proactively selecting the salient visual regions.
@@ A.2 Baselines
A.3 Template of $\mathcal{P}_{CAG}$
Figure 3 provides an intuitive example showing the template of $\mathcal{P}_{CAG}$ and how it is used to prompt the VLM to carefully generate a guiding description for the input image. In particular, for multiple-choice questions, such as those in M3CoT and ScienceQA, we prepend the following brief explanation to $\mathcal{P}_{CAG}$ to aid the VLM in better understanding its designated task: “This is a multiple-choice question. The question is based on the image provided.
@@ A.3 Template of $\mathcal{P}_{CAG}$
A.4 Setting of Hyper-parameters
The hyper-parameter settings are listed in Table 7.